# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ VI Agent-Agent Communication
| CS | Agent Communication Protocol | IBM | ACP-IBM[119] | Four agent discovery mechanisms, synchronous and streaming execution, multi-turn state preservation |
| Agent Connect Protocol | AGNTCY | ACP-AGNTCY[47] | Allow authenticating callers, threaded state management, flexible execution model | |
| Agent Communication Protocol | AgentUnion | ACP-AgentUnion[6] | Decentralized APs based on the existing domain name system, each AP holds its agent list | |
| P2P | Agent Communication Network | Fetch.AI | ACN[230] | Distributed-Hash-Table-based peer-to-peer discovery, end-to-end encryption. |
| Agent Network Protocol | ANP Team | ANP[263] | A three-layer architecture and W3C-compliant Decentralized Identifiers. | |
| Layered Orchestration for Knowledgeful Agents | CMU | LOKA[231] | Decentralized identifier, intent-centric communication, privacy-preserving accountability, ethical governance | |
| Hybrid | Language Model Operating System Protocol | Eclipse | LMOS[76] | Three agent discovery mechanisms, decentralized digital identifiers, dynamic transport protocol support, group management. |
| Agent to Agent Protocol | A2A[89] | Three agent discovery mechanisms, cross-platform compatibility, asynchronous priority, security mechanisms | ||
| Others | Agora | Oxford | Agora[198] | Dynamically switches communication modes based on the communication frequency |
| Agent Protocol | LangChain | Agent Protocol[153] | Flexible communication mechanisms based on Run, Thread, and Store. | |
| Agent Interaction & Transaction Protocol | NEAR AI | AITP[9] | Threads-based communication, secure communication across trust boundaries. |
VI-A Protocols
@@ VI-A1 CS-based Communication
As shown in Figure 7, CS-based communication protocols follow the client-server architecture, which provides centralized servers to manage the information of agents (e.g., their unique IDs and capability descriptions). Under this paradigm, agents interact through well-defined interfaces and rely on centralized servers to discover the desired agents. CS-based communication offers stronger agent discovery functionality, such as supporting the search of agents based on capabilities. For example, the agent servers can run complex search/match algorithms to find proper agent descriptions in their databases.
@@
ACP-IBM. The Agent Communication Protocol proposed by IBM is designed for the collaboration of agents [119]. We call it ACP-IBM in this paper to distinguish it from the Agent Communication Protocols proposed by other organizations. In ACP-IBM, the client is connected to an agent server. First, the client conducts an agent discovery process to discover available agents and get the description of their capabilities. ACP-IBM supports different discovery mechanisms such as Basic Discovery, Registry-Based Discovery, Offline Discovery, and Open Discovery. Second, after confirming the agent(s), the client starts the invocation. As shown in Figure 8, for a single-agent task, the agent server wraps the agent, translating REST calls into internal logic. For multi-agent tasks, the client message is first sent to a Router Agent, which is responsible for decomposing requests, routing tasks, and aggregating responses. ACP-IBM supports synchronous and streaming execution, and allows the preservation of state across multi-turn conversations.
@@
ACP-AGNTCY. The Agent Connect Protocol proposed by AGNTCY [47] is an open standard designed to facilitate seamless communication between agents. The client can first search available agents on the agent server, which returns a list of agent IDs matching the criteria provided in the request. Then, the client further gets the agent descriptor by agent ID to know the detailed description of the agent functionality. After confirming the target agent, the client can assign tasks to this agent and wait for results. The characteristics of ACP-AGNTCY include flexibility and scalability. First, ACP-AGNTCY deploys a Threads Mechanism, which enables contextual continuity, supporting the creation, copying, and searching of threads, and recording state histories for debugging and backtracking. Second, it supports two operation modes: stateless and stateful. The former is suitable for simple single tasks, while the latter supports multi-round conversations, state continuation, and historical data traceability through the thread mechanism to meet the requirements of complex scenarios.
@@
ACP-AgentUnion. The Agent Communication Protocol proposed by AgentUnion [6] also aims to provide seamless communication among heterogeneous agents. Each agent has a unique AID (Agent ID), which is a secondary domain name (i.e., agent_name.ap_domain). Agents access IoA through the AP (Access Point), which completes the agent’s identity authentication, address search, communication, and data storage, and provides AID creation, management, and authentication services. As a result, APs can provide the proper agent list based on the query from users. In this way, agents can communicate with other agents on the Internet.
VI-A2 Peer-to-Peer-based Communication
@@
As shown in Figure 7, P2P-based communication protocols pursue a decentralized agent discovery mechanism. They usually wish to use globally universal identifiers (e.g., combined with a domain name) to enable agents to directly search other agents on the Internet. The advantage of this paradigm is that it supports convenient location and global search (e.g., using a crawler) of agents, but they usually do not support the discovery based on agent capability.
@@
ACN. Agent Communication Network (ACN) [230] is a decentralized, peer-to-peer communication infrastructure to facilitate secure and efficient interactions among agents without relying on centralized coordination. Leveraging a Distributed Hash Table (DHT), ACN enables agents to publish and discover public keys, allowing for the establishment of encrypted, point-to-point communication channels. First, agents need to register with one peer node, and the peer node stores the “agent ID - peer node ID” pair in the DHT network. Then, during communication, the source agent sends the message to its associated peer node, and this node recursively searches for the peer node of the target agent through DHT: if the target record exists, the peer nodes of both parties establish a direct communication channel, and forward the message after digital signature verification; if not, an error is returned. The entire communication process uses end-to-end encryption (e.g., TLS) to ensure security. Like the Well-Known URI discovery of A2A, ACN does not support the discovery based on agent capabilities.
@@
ANP (Agent Network Protocol) [263] is an open communication framework designed to enable scalable and secure interoperability among heterogeneous autonomous agents. It supports two types of agent discovery: active and passive. The active discovery uses a uniform URI (.well-known), while the passive discovery submits the agent description to search services. ANP employs a three-layer architecture. At the Identity and Encrypted Communication Layer, it leverages W3C-compliant Decentralized Identifiers (DIDs) and end-to-end Elliptic Curve Cryptography (ECC) encryption to ensure verifiable cross-platform authentication and confidential agent communication. The Meta-Protocol layer allows agents to dynamically establish and evolve communication protocols through natural language interaction, enabling flexible, adaptive, and efficient inter-agent coordination. At the Application layer, ANP describes agent capabilities using JSON-LD and semantic web standards such as RDF and schema.org, enabling agents to discover and invoke services based on semantic descriptions. It also defines standardized protocol management mechanisms to support efficient and interoperable agent interaction. From a security standpoint, ANP enforces the separation of human authorization from agent-level delegation and adheres to the principle of least privilege. Its minimal-trust, modular design aims to eliminate platform silos and foster a decentralized, composable agent ecosystem.
@@
LOKA. LOKA (Layered Orchestration for Knowledgeful Agents) protocol [231] aims to build a trustworthy and ethical agent ecosystem. Its principle is based on the collaborative operation of multiple key components. First, LOKA introduces the Universal Agent Identity Layer (UAIL), using Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs) to assign each agent a unique and verifiable identity, thereby achieving decentralized identity management and autonomous control. Second, LOKA proposes an Intent-Centric Communication Protocol, which supports the exchange of semantically rich and ethically annotated messages among agents, promoting semantic coordination and efficient communication. Third, LOKA proposes the Decentralized Ethical Consensus Protocol (DECP). DECP uses multi-party computation (MPC) and distributed ledger technology to enable agents to make context-aware decisions based on a shared ethical baseline, ensuring that their behavior complies with ethical norms. In addition, the authors also point out that it combines cutting-edge technologies such as distributed identity, verifiable credentials, and post-quantum cryptography to provide comprehensive support for the agent ecosystem in terms of identity management, communication and coordination, ethical decision-making, and security.
VI-A3 Hybrid Communication
@@
Hybrid communication protocols support both CS-based and P2P-based agent discovery. However, please note that such support is determined by different scenarios. For example, they usually propose a CS-based discovery mechanism specifically for local area networks, while the worldwide agent discovery is still P2P-based. In other words, although such protocols support more flexible agent discovery to fit different scenarios, they do not completely eliminate the existing limitations of agent discovery.
@@
LMOS. The LMOS (Language Model Operating System) Protocol proposed by Eclipse [76] aims to enable agents and tools from diverse organizations to be easily discovered and connected, regardless of the technologies they are built on. LMOS supports three different agent discovery methods to enable both centralized and decentralized discovery. The first method is to adopt the mechanism of W3C Web of Things (WoT) to enable agents to dynamically register metadata in the registry. The second method is to use mDNS and DNS-SD protocol to discover agents/tools in local area networks. The last method is adopting a federal, decentralized protocol (such as P2P protocol) to disseminate agents and tool descriptions, without relying on a centralized registry center, which is applicable for global collaboration of agents. The LMOS protocol has a three-layer architecture. The Application Layer utilizes a JSON-LD-based format to describe the capabilities of agents and tools. The Transport Layer facilitates flexible communication by enabling agents to negotiate protocols like HTTP or MQTT dynamically, accommodating both synchronous and asynchronous data exchange to suit different use cases. The Identity and Security Layer establishes trust through W3C-compliant decentralized identity authentication, combined with encryption and protocols like OAuth2, to secure cross-platform interactions.
@@
A2A. The A2A (Agent to Agent) Protocol proposed by Google [89] aims to enable collaboration between agents. A2A supports three different mechanisms for agent discovery. The first is Well-Known URI, which requires agent servers to store Agent Cards in standardized “well-known” paths under the domain name (e.g., https://agent-server-domain/.well-known/agent.json). This mechanism enables automatic search of agents on the Internet. However, it does not support the discovery of agents based on capabilities. The second is Curated Registries, i.e., agent servers register their Agent Cards, which is similar to ACP-IBM. The above two methods can be referred to as Figure 7. The third is Direct Configuration / Private Discovery, which means that the client can directly require Agent Cards through hard-coded, local configuration files, environment variables, or private APIs. After finding the desired agents, the client can assign tasks to them and wait for the responses.
VI-A4 Others
@@
This kind of protocols do not explicitly illustrate their unique design for agent discovery. Instead, they only focus on the communication process, e.g., the data format, the management of multiple queries, or the historical conversation state.
@@
Agora. Agora [198] is a communication protocol for the communication of heterogeneous agents. Its core mechanism dynamically switches communication modes based on the communication frequency. Specifically, standardized protocols manually developed (such as OpenAPI) are used for high-frequency communications to ensure efficiency. Natural language processed by agents is employed for low-frequency or unknown scenarios to maintain versatility. Structured data handled by the routines (written by agents) is utilized for intermediate-frequency communications to balance cost and flexibility. Meanwhile, Protocol Documents (PDs) are used as self-contained protocol descriptions, uniquely identified by hash values and supporting decentralized sharing, enabling agents to autonomously negotiate and reuse protocols without a central authority. In the Agora network, there are multiple protocol databases that store PDs. Each Agent can submit the negotiated protocol documents to the database for other Agents to retrieve and use. These databases use peer-to-peer synchronization: different protocol databases will share protocol documents regularly (e.g., after every 10 queries), enabling cross-database dissemination of protocols. Agora is also compatible with existing communication standards, allowing agents to independently develop and share protocols during communication, achieving automated processing of complex tasks in large-scale networks.
@@
AITP (Agent Interaction & Transaction Protocol) [9] is a standardized framework that enables structured and interoperable communication among agents. AITP deploys a thread-based messaging structure. Each thread encapsulates the conversational context, participant metadata, and capability declarations, supporting consistent multi-agent coordination across heterogeneous environments. The protocol employs JSON-formatted message exchanges to encode requests, responses, and contextual information. It supports both synchronous and asynchronous interaction patterns, facilitating the orchestration of complex, multi-step tasks. AITP does not provide specific agent discovery mechanisms. It focuses on the communication process of agents.
@@
Agent Protocol. Agent Protocol is proposed by LangChain [153] to enable the communication between LanghGraph (a multi-agent framework) and other types of agents. Its mechanism is based on Thread and Run: Run is a single call of the agent, which supports streaming output of real-time results or waiting for the final output. Threads act as state containers. They store the cumulative output and checkpoints of multiple rounds of operation. Besides, they support the management of state history (such as querying, copying, and deleting), ensuring that the agent maintains context continuity during multiple rounds of calls. Furthermore, Background Runs support asynchronous task processing, and progress can be managed through an independent interface. The element Store provides cross-thread persistent key-value storage for achieving long-term memory. The overall mechanism realizes flexible control over proxy calls, status management, asynchronous tasks and data storage through HTTP interfaces and configuration parameters. Agent Protocol does not explicitly illustrate the unique agent discovery mechanism it supports.
VI-B Security Risk Analysis

### v4

@@ VI Agent-Agent Communication
| CS | Agent Communication Protocol | IBM | ACP-IBM[ibm_acp_2024] | Four agent discovery mechanisms, synchronous and streaming execution, multi-turn state preservation |
| Agent Connect Protocol | AGNTCY | ACP-AGNTCY[langchain_aconp] | Allow authenticating callers, threaded state management, flexible execution model | |
| Agent Communication Protocol | AgentUnion | ACP-AgentUnion[ACPagentunion] | Decentralized APs based on the existing domain name system, each AP holds its agent list | |
| P2P | Agent Communication Network | Fetch.AI | ACN[p2pacn] | Distributed-Hash-Table-based peer-to-peer discovery, end-to-end encryption. |
| Agent Network Protocol | ANP Team | ANP[anp] | A three-layer architecture and W3C-compliant Decentralized Identifiers. | |
| Layered Orchestration for Knowledgeful Agents | CMU | LOKA[ranjan2025loka] | Decentralized identifier, intent-centric communication, privacy-preserving accountability, ethical governance | |
| Hybrid | Language Model Operating System Protocol | Eclipse | LMOS[eclipse_lmos] | Three agent discovery mechanisms, decentralized digital identifiers, dynamic transport protocol support, group management. |
| Agent to Agent Protocol | A2A[A2A] | Three agent discovery mechanisms, cross-platform compatibility, asynchronous priority, security mechanisms | ||
| Others | Agora | Oxford | Agora[marro2024scalable] | Dynamically switches communication modes based on the communication frequency |
| Agent Protocol | LangChain | Agent Protocol[AGENTprotocollangchain] | Flexible communication mechanisms based on Run, Thread, and Store. | |
| Agent Interaction & Transaction Protocol | NEAR AI | AITP[nearai_aitp] | Threads-based communication, secure communication across trust boundaries. |
VI-A Communication Protocols
As shown in Figure 5, the communication between agents may be in a WAN or a LAN. For the former, agents need secure protocols such as HTTPS, which is the same as user-agent interaction. Besides, agents may communicate in a LAN, which has a more secure network environment. In this context, to reduce communication overhead and improve efficiency, agents can directly communicate in plaintext, such as using HTTP or customized protocols on IP packets. Due to the same reason in the user-agent interaction (Section V-A), protocols for agent-agent communication are also about L2.
@@ VI-A Communication Protocols
(i) CS-based Communication. As shown in Figure 11, CS-based communication protocols follow the client-server architecture, which provides centralized servers to manage the information of agents (e.g., their unique IDs and capability descriptions). Under this paradigm, agents interact through well-defined interfaces and rely on centralized servers to discover the desired agents. CS-based communication offers stronger agent discovery functionality, such as supporting the search for agents based on capabilities. For example, the agent servers can run complex search/match algorithms to find proper agent descriptions in their databases.
@@
ACP-IBM. The Agent Communication Protocol proposed by IBM is designed for the collaboration of agents [ibm_acp_2024]. We call it ACP-IBM in this paper to distinguish it from the Agent Communication Protocols proposed by other organizations. In ACP-IBM, the client is connected to an agent server. First, the client conducts an agent discovery process to discover available agents and get the description of their capabilities. ACP-IBM supports different discovery mechanisms such as Basic Discovery, Registry-Based Discovery, Offline Discovery, and Open Discovery. Second, after confirming the agent(s), the client starts the invocation. As shown in Figure 12, for a single-agent task, the agent server wraps the agent, translating REST calls into internal logic. For multi-agent tasks, the client message is first sent to a Router Agent, which is responsible for decomposing requests, routing tasks, and aggregating responses. ACP-IBM supports synchronous and streaming execution, and allows the preservation of state across multi-turn conversations.
@@
ACP-AGNTCY. The Agent Connect Protocol proposed by AGNTCY [langchain_aconp] is an open standard designed to facilitate seamless communication between agents. The client can first search available agents on the agent server, which returns a list of agent IDs matching the criteria provided in the request. Then, the client further gets the agent descriptor by agent ID to know the detailed description of the agent functionality. After confirming the target agent, the client can assign tasks to this agent and wait for results. The characteristics of ACP-AGNTCY include flexibility and scalability. First, ACP-AGNTCY deploys a Threads Mechanism, which enables contextual continuity, supporting the creation, copying, and searching of threads, and recording state histories for debugging and backtracking. Second, it supports two operation modes: stateless and stateful. The former is suitable for simple single tasks, while the latter supports multi-round conversations, state continuation, and historical data traceability through the thread mechanism to meet the requirements of complex scenarios.
@@
ACP-AgentUnion. The Agent Communication Protocol proposed by AgentUnion [ACPagentunion] also aims to provide seamless communication among heterogeneous agents. Each agent has a unique AID (Agent ID), which is a secondary domain name (i.e., agent_name.ap_domain). Agents access IoA through the AP (Access Point), which completes the agent’s identity authentication, address search, communication, and data storage, and provides AID creation, management, and authentication services. As a result, APs can provide the proper agent list based on the query from users. In this way, agents can communicate with other agents on the Internet.
@@
(ii) Peer-to-Peer-based Communication. As shown in Figure 11, P2P-based communication protocols pursue a decentralized agent discovery mechanism. They usually wish to use globally universal identifiers (e.g., combined with a domain name) to enable agents to directly search other agents on the Internet. The advantage of this paradigm is that it supports convenient location and global search (e.g., using a crawler) of agents, but they usually do not support the discovery based on agent capability.
@@
ACN. Agent Communication Network (ACN) [p2pacn] is a decentralized, peer-to-peer communication infrastructure to facilitate secure and efficient interactions among agents without relying on centralized coordination. Leveraging a Distributed Hash Table (DHT), ACN enables agents to publish and discover public keys, allowing for the establishment of encrypted, point-to-point communication channels. First, agents need to register with one peer node, and the peer node stores the “agent ID - peer node ID” pair in the DHT network. Then, during communication, the source agent sends the message to its associated peer node, and this node recursively searches for the peer node of the target agent through DHT: if the target record exists, the peer nodes of both parties establish a direct communication channel and forward the message after digital signature verification; if not, an error is returned. The entire communication process uses end-to-end encryption (e.g., TLS) to ensure security. Like the Well-Known URI discovery of A2A, ACN does not support the discovery based on agent capabilities.
@@
ANP. Agent Network Protocol (ANP) [anp] is an open communication framework designed to enable scalable and secure interoperability among heterogeneous autonomous agents. It supports two types of agent discovery: active and passive. The active discovery uses a uniform URI (.well-known), while the passive discovery submits the agent description to search services. ANP employs a three-layer architecture. At the Identity and Encrypted Communication Layer, it leverages W3C-compliant Decentralized Identifiers (DIDs) and end-to-end Elliptic Curve Cryptography (ECC) encryption to ensure verifiable cross-platform authentication and confidential agent communication. The Meta-Protocol layer allows agents to dynamically establish and evolve communication protocols through natural language interaction, enabling flexible, adaptive, and efficient inter-agent coordination. At the Application layer, ANP describes agent capabilities using JSON-LD and semantic web standards such as RDF and schema.org, enabling agents to discover and invoke services based on semantic descriptions. It also defines standardized protocol management mechanisms to support efficient and interoperable agent interaction. From a security standpoint, ANP enforces the separation of human authorization from agent-level delegation and adheres to the principle of least privilege. Its minimal-trust, modular design aims to eliminate platform silos and foster a decentralized, composable agent ecosystem.
@@
LOKA. Layered Orchestration for Knowledgeful Agents (LOKA) protocol [ranjan2025loka] aims to build a trustworthy and ethical agent ecosystem. Its principle is based on the collaborative operation of multiple key components. First, LOKA introduces the Universal Agent Identity Layer (UAIL), using Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs) to assign each agent a unique and verifiable identity, thereby achieving decentralized identity management and autonomous control. Second, LOKA proposes an Intent-Centric Communication Protocol, which supports the exchange of semantically rich and ethically annotated messages among agents, promoting semantic coordination and efficient communication. Third, LOKA proposes the Decentralized Ethical Consensus Protocol (DECP). DECP uses multi-party computation (MPC) and distributed ledger technology to enable agents to make context-aware decisions based on a shared ethical baseline, ensuring that their behavior complies with ethical norms. In addition, the authors also point out that it combines cutting-edge technologies such as distributed identity, verifiable credentials, and post-quantum cryptography to provide comprehensive support for the agent ecosystem in terms of identity management, communication and coordination, ethical decision-making, and security.
@@
(iii) Hybrid Communication. Hybrid communication protocols support both CS-based and P2P-based agent discovery. However, please note that such support is determined by different scenarios. For example, they usually propose a CS-based discovery mechanism specifically for local area networks, while the worldwide agent discovery is still P2P-based. In other words, although such protocols support more flexible agent discovery to fit different scenarios, they do not completely eliminate the existing limitations of agent discovery.
@@
LMOS. The Language Model Operating System (LMOS) Protocol proposed by Eclipse [eclipse_lmos] aims to enable agents and tools from diverse organizations to be easily discovered and connected, regardless of the technologies they are built on. LMOS supports three different agent discovery methods to enable both centralized and decentralized discovery. The first method is to adopt the mechanism of W3C Web of Things (WoT) to enable agents to dynamically register metadata in the registry. The second method is to use mDNS and the DNS-SD protocol to discover agents/tools in local area networks. The last method is adopting a federal, decentralized protocol (such as a P2P protocol) to disseminate agents and tool descriptions, without relying on a centralized registry center, which is applicable for global collaboration of agents. The LMOS protocol has a three-layer architecture. The Application Layer utilizes a JSON-LD-based format to describe the capabilities of agents and tools. The Transport Layer facilitates flexible communication by enabling agents to negotiate protocols like HTTP or MQTT dynamically, accommodating both synchronous and asynchronous data exchange to suit different use cases. The Identity and Security Layer establishes trust through W3C-compliant decentralized identity authentication, combined with encryption and protocols like OAuth2, to secure cross-platform interactions.
@@
A2A. The Agent to Agent (A2A) Protocol proposed by Google [A2A] aims to enable collaboration between agents. A2A supports three different mechanisms for agent discovery. The first is Well-Known URI, which requires agent servers to store Agent Cards in standardized “well-known” paths under the domain name (e.g., https://agent-server-domain/.well-known/agent.json). This mechanism enables automatic search of agents on the Internet. However, it does not support the discovery of agents based on capabilities. The second is Curated Registries, i.e., agent servers register their Agent Cards, which is similar to ACP-IBM. The above two methods can be referred to as Figure 11. The third is Direct Configuration / Private Discovery, which means that the client can directly require Agent Cards through hard-coded, local configuration files, environment variables, or private APIs. After finding the desired agents, the client can assign tasks to them and wait for the responses.
@@
(iv) Others. These kinds of protocols do not explicitly illustrate their unique design for agent discovery. Instead, they only focus on the communication process, e.g., the data format, the management of multiple queries, or the historical conversation state.
@@
Agora. Agora [marro2024scalable] is a communication protocol for the communication of heterogeneous agents. Its core mechanism dynamically switches communication modes based on the communication frequency. Specifically, standardized protocols manually developed (such as OpenAPI) are used for high-frequency communications to ensure efficiency. Natural language processed by agents is employed for low-frequency or unknown scenarios to maintain versatility. Structured data handled by the routines (written by agents) is utilized for intermediate-frequency communications to balance cost and flexibility. Meanwhile, Protocol Documents (PDs) are used as self-contained protocol descriptions, uniquely identified by hash values and supporting decentralized sharing, enabling agents to autonomously negotiate and reuse protocols without a central authority. In the Agora network, there are multiple protocol databases that store PDs. Each Agent can submit the negotiated protocol documents to the database for other Agents to retrieve and use. These databases use peer-to-peer synchronization: different protocol databases will share protocol documents regularly (e.g., after every 10 queries), enabling cross-database dissemination of protocols. Agora is also compatible with existing communication standards, allowing agents to independently develop and share protocols during communication, achieving automated processing of complex tasks in large-scale networks.
@@
AITP. Agent Interaction & Transaction Protocol [nearai_aitp] is a standardized framework that enables structured and interoperable communication among agents. AITP deploys a thread-based messaging structure. Each thread encapsulates the conversational context, participant metadata, and capability declarations, supporting consistent multi-agent coordination across heterogeneous environments. The protocol employs JSON-formatted message exchanges to encode requests, responses, and contextual information. It supports both synchronous and asynchronous interaction patterns, facilitating the orchestration of complex, multi-step tasks. AITP does not provide specific agent discovery mechanisms. It focuses on the communication process of agents.
@@
Agent Protocol. Agent Protocol is proposed by LangChain [AGENTprotocollangchain] to enable the communication between LanghGraph (a multi-agent framework) and other types of agents. Its mechanism is based on Thread and Run: Run is a single call of the agent, which supports streaming output of real-time results or waiting for the final output. Threads act as state containers. They store the cumulative output and checkpoints of multiple rounds of operation. Besides, they support the management of state history (such as querying, copying, and deleting), ensuring that the agent maintains context continuity during multiple rounds of calls. Furthermore, Background Runs support asynchronous task processing, and progress can be managed through an independent interface. The element Store provides cross-thread persistent key-value storage for achieving long-term memory. The overall mechanism realizes flexible control over proxy calls, status management, asynchronous tasks, and data storage through HTTP interfaces and configuration parameters. Agent Protocol does not explicitly illustrate the unique agent discovery mechanism it supports.
VI-B Security Risks